OpenStudy (anonymous):

Find the length and the midpoint of the segment joining the two given points. K (23,4) and W (-2,16)

OpenStudy (anonymous):

does it have a graph or something?

OpenStudy (anonymous):

a. 30; (0, 0) b. 27; (10.5, 10) c. 15; (4, 9) d. 45; (-8, 9)

OpenStudy (anonymous):

if you want to find the midpoint of these points use these formula\[ x_{m}=\frac{ x _{k}+x _{w} }{ ? }\]\[y _{m}=\frac{ y _{k}+y _{w} }{ 2 }\]

OpenStudy (anonymous):

i don't know how, do you know the answer?

OpenStudy (anonymous):

\[x _{m}=\frac{ 23-2 }{ 2 }=\frac{ 21 }{ 2 }=10.5\]\[y _{m}=\frac{ 4+16 }{ 2 }=\frac{ 20 }{ 2 }=10\] so m=(10.5,10)

OpenStudy (anonymous):

B is true answer.

OpenStudy (anonymous):

(10, –8), (4, 0) a. 2 b. 10 c. 16 d. 100

OpenStudy (anonymous):

the distance between the two points?